OpenStudy (anonymous):

Please go through with me how to solve this problem.. You have three $1 bills, four $5 bills, and two $10 bills in your wallet. You select a bill at random. Without replacing the bill, you choose a second bill at random. Find P($10 then $1). A. 1/12 B. 5/81 C. 5/72 D. 2/27

OpenStudy (anonymous):

Could you help me with this? @agent0smith

OpenStudy (mertsj):

What is the probability that will you get $10 on the first draw?

OpenStudy (anonymous):

Well $10 out of $43 ... would I just make that a fraction? Or .... I don't even know >.<

OpenStudy (mertsj):

There are 9 bills in your wallet. 2 of them are tens. The probability you will get a 10 is 2/9

OpenStudy (mertsj):

So you draw out the 10. How many bills are in your wallet now?

OpenStudy (anonymous):

If you draw out one of the tens, then there would be 8 bills left, both tens, would be 7..

OpenStudy (mertsj):

You had 9 bills in your wallet. You took out 1 bill. 9-1=8. You now have 8 bills in your wallet. How many of them are $1 bills?

OpenStudy (anonymous):

3 of them are $! bills

OpenStudy (anonymous):

$1 bills, lol my fault

OpenStudy (mertsj):

So what is the probability you will draw out a $1 bill on your next draw?

OpenStudy (anonymous):

3/8 ?

OpenStudy (mertsj):

Yes. So now you should multiply those individual probabilities to get your answer: \[\frac{2}{9}\times\frac{3}{8}\]

OpenStudy (anonymous):

Do i cross multiply or multiply the numerator x numerator and denominator x denominator?

OpenStudy (mertsj):

Cancel the two and then multiply numerators and multiply denominators.

OpenStudy (anonymous):

What do you mean by cancel the two? o.o

OpenStudy (mertsj):